Circular No. 6064

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

RX J0045.4+4154 N. E. White, Laboratory for High Energy Astrophysics, Goddard Space

Flight Center (GSFC); P. Giommi, European Space Information System, European Space Agency; L. Angelini, GSFC; and S. Fantasia, University of Maryland, report the discovery of a recurrent supersoft x-ray transient in ROSAT archival data on M31: “The first outburst began on 1992 Feb. 2 and lasted at least four days (the end of that particular ROSAT observation sequence). The source had ‘turned off’ by the time of the next ROSAT observation of this region, which began on 1992 Aug. 10. A second outburst began on 1993 Jan. 7; the peak of the outburst was reached in ∼ 4 days, after which a decline started. When the ROSAT observation ended 2.5 days after the peak, the source had not yet reached the minimum. Fitting of the source spectra on both occasions yields a blackbody temperature of 0.095 keV, consistent with the definition of a supersoft source. The hydrogen column density is 1.2 × 10^{21} cm^{−2} and the unabsorbed source luminosity in the band 0.1–4.0 keV is 3.4 × 10^{38} erg/s (assuming the source is located in M31). The source position is α = 0^{h}45^{m}28^{s}.6, δ = +41^{o}54′11′′.3 (equinox 2000.0; 90-percent confidence radius 13′′). In the ROSAT archive there are four time periods in which this part of the sky was observed, each separated by six months (1991 July, 1992 Jan.–Feb. and Aug., 1993 Jan.). On two occasions each the source was in outburst (1992 Jan.–Feb. and 1993 Jan.) and in quiescence. If this source is not a foreground object, then it is the first recurrent x-ray transient to be found in M31. Further outbursts might be expected. If they are periodic, the next will occur in 1994 Nov. Further observations are suggested.”

PSR 0833–45 C. Flanagan, Hartebeesthoek Radio Astronomy Observatory, Johan-

nesburg, writes: “PSR 0833–45 (the Vela pulsar) has undergone a further spin-up, the eleventh observed to date, 32 days after the previous such event (IAUC 6038). The most recent event occurred on Aug. 27.158, and involved a relative change in spin-period dP/P of −1.99(2) × 10^{−7}, 23 percent the size of the glitch of July 26. Unlike the July 26 event, for which the observed relative change in period derivative < 0.005, the Aug. 27 event was accompanied by an increase in period derivative dP′/P′ of +0.12(2).”

1994 August 29 Daniel W. E. Green